OpenStudy (thatcluelessgirl):

what is 4550kilometers divided by 2.5cm

OpenStudy (thekingofcr):

182 000 000 kilo

OpenStudy (thatcluelessgirl):

what is that in cm and meters

OpenStudy (royalranger):

Depends on if they want it in meters or cm

OpenStudy (thatcluelessgirl):

ok hold on

OpenStudy (thekingofcr):

1.82e+13 cm 182000000000 meter

OpenStudy (thatcluelessgirl):

and then 100 cm in 1 meter

OpenStudy (thekingofcr):

18200000000000 cm 182000000000 meter

OpenStudy (thekingofcr):

got it?

OpenStudy (thatcluelessgirl):

thank you yes hun I did

OpenStudy (thekingofcr):

you did what?

OpenStudy (thatcluelessgirl):

I got it

OpenStudy (thekingofcr):

ok

OpenStudy (sshayer):

\[\frac{ 4550 \times 1000 \times 100 }{ 2.5 }=?\] 1 km =1000 m 1 m=100 cm